# 2024_subroc.bib

### v1

@@ Abstract
Machine learning (ML) is increasingly employed in real-world applications like medicine or economics, thus, potentially affecting large populations. However, ML models often do not perform homogeneously across such populations resulting in subgroups of the population (e.g., sex=female $\wedge$ marital_status=married) where the model underperforms or, conversely, is particularly accurate. Identifying and describing such subgroups can support practical decisions on which subpopulation a model is safe to deploy or where more training data is required. The potential of identifying and analyzing such subgroups has been recognized, however, an efficient and coherent framework for effective search is missing. Consequently, we introduce SubROC, an open-source, easy-to-use framework based on Exceptional Model Mining for reliably and efficiently finding strengths and weaknesses of classification models in the form of interpretable population subgroups. SubROC incorporates common evaluation measures (ROC and PR AUC), efficient search space pruning for fast exhaustive subgroup search, control for class imbalance, adjustment for redundant patterns, and significance testing. We illustrate the practical benefits of SubROC in case studies as well as in comparative analyses across multiple datasets.
@@ 1 Introduction
Machine learning (ML) is widely adopted in high-stake application domains such as health care, loans, hiring decisions, or legal decision-making [mehrabi_survey_2021]. Thus, ML models and their decisions already influence large populations and everyday life. However, it has also repeatedly been shown that ML models can be highly heterogeneous in their performance [mehrabi_survey_2021]. Subgroups of a population may exist where a ML model underperforms (or performs remarkably well). For illustration, see Figure 1 where predictions are particularly bad for the subpopulation of women from rural areas (sex=female $\wedge$ housing=rural) and remarkably accurate for male teenagers (sex=male $\wedge$ $\leq$18). Knowing such subgroups may allow data scientists and practitioners to make informed decisions about model deployment and development. Here, the practitioners may decide to use the current model only for male teenagers, where the model performs well, and the data scientists may decide to collect more data for subgroups where the model underperforms.
@@
Figure 1: In an exemplary application, binary classification could distinguish between healthy and diseased patients (left). Predictions of soft classifiers impose an ordering on the persons, which performs better the clearer healthy patients are separated from diseased patients. In this case there is a clear difference between the densities of healthy and diseased patients in this ordering, which are represented by the red and blue curves here (middle left). A dataset can be split up in different ways by patterns describing subsets. Our approach based on subgroup discovery evaluates these subsets systematically and returns those where the model predictions create the best or worst separation of the classes (middle right). Results may highlight arbitrary sensitive subgroups of a population, where the model under- or over-performs compared to the rest of the dataset (right).
@@
Background and problem setting. The need and potential of identifying and analyzing subgroups has been widely recognized. For example, a wide array of work aims at alleviating consistent underperformance in known [sagawa_distributionally_2020] as well as previously unknown [sohoni_no_2020] subgroups. Similar observations have been made concerning the fairness of ML models [mehrabi_survey_2021]. Consequently, various approaches have been explored to study such atypical subgroups systematically, e.g., with interactive tools for manual discovery of subgroups [baylor_tfx_2017, kahng_visual_2016]. However, this is labor-intensive, inefficient, and poses the threat of missing essential subgroups. Thus, automated approaches have been proposed that enable the search for interpretable subgroups where ML models over- or underperform [chung_automated_2020, cabrera_fairvis_2019]. The most principled approach employs the framework of Exceptional Model Mining (EMM) with a custom performance measure for efficient exhaustive search [duivesteijn_understanding_2014]. However, none of these approaches combine the automated and exhaustive search of EMM, standard measures for soft classifiers like ROC AUC or PR AUC, efficient search, and handling of class imbalance and statistical significance testing, which are required for a routine and practical analysis.
@@
Contribution. Consequently, we introduce SubROC, which builds on EMM to reliably and efficiently find strengths and weaknesses of classification models. For this, we (a) propose a new group of scoring functions based on the area-under-the-curve (AUC) performance statistic, (b) derive tight optimistic estimates that enable fast exhaustive subgroup search, (c) provide a well-rounded framework that includes control for class imbalance, pruning of redundant results, and significance testing, (d) show the practical benefits of our approach across multiple datasets, and (e) integrate SubROC in the easy-to-use, open source pysubgroup package and provide the code for our experiments.  [^1] Experiment code is available at https://github.com/bckrlab/subroc.
@@
Overall, SubROC will enable practitioners to routinely evaluate their ML models by finding subgroups where they over- or underperform. This will not only have the potential to impact and improve quality control procedures and increase the confidence in deployed ML models but also lay the foundation for downstream approaches to systematically improve their performance.
@@ 2.1 Soft Classifier Evaluation
SubROC focuses on binary soft classification. $D=(\mathcal{I},\mathcal{A})$ is a dataset with instances $c\in\mathcal{I}$ and attributes (features) $\mathcal{A}=(X_{1},\ldots,X_{m})$. $X_{i}(c)$ denotes the value of attribute $X_{i}$ in instance $c$. $Y(c)\in\{1,0\}$ is the ground truth label and $\hat{Y}(c)\in[0,1]$ a continuous (soft) prediction. Based on the ground truth labels, we call some instances positives ($\mathcal{P}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=1\}$) and some negatives ($\mathcal{N}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=0\}$). A classifier performs well if it returns high (low) values for instances with the ground truth $Y(c)=1$ ($Y(c)=0$).
@@
While many evaluation metrics for soft classification exist, including binary cross-entropy or the Brier score, the area under the curve (AUC) of the Receiver Operating Characteristic (ROC) curve is among the most widely used. It assumes a ranking of instances according to their predicted probabilities from high to low. Then, it computes the probability of a randomly chosen positive instance being ranked above a randomly chosen negative instance.
@@
A main reason for the popularity of AUC-based evaluation metrics is that they summarize across all potential thresholds that a user could apply to conduct a hard classification into two classes based on the soft classifier results, e.g., depending on misclassification costs. This makes the ROC AUC a natural choice for interpretable subgroup descriptions, contributing to the insights that subgroups from our framework may provide to experts.
2.2 Subgroup Discovery and Exceptional Model Mining
@@
A selector $s\in\mathcal{S}$ is a boolean function based on a single attribute, e.g., $sex=female$ or $age\in[20,30)$. A pattern $p\in 2^{\mathcal{S}}$ is a conjunction of selectors, e.g., $sex=female\wedge age\in[20,30)$. In a given dataset, a pattern’s cover $sg(p)$ represents the set of all instances covered by it, i.e., all instances with $p=true$. A pattern and its cover are referred to as a subgroup. We call a subgroup $p_{spec}$ a specialization of another subgroup $p_{gen}$ (its generalization) if its pattern’s conjunction of selectors adds at least one additional selector: $p_{gen}\subset p_{spec}$.
@@
Subgroup discovery (SD) [klosgen_explora_1996] finds subgroups that show an interesting deviation with respect to a target concept, which is traditionally given as a single data attribute. E.g., an interesting subgroup may contain a higher-than-average disease occurrence. EMM [leman_exceptional_2008] extends target concepts to model classes. Here, a subgroup is considered interesting if the model parameters fitted to the subgroup are significantly different from those fitted to the complete dataset.
@@
SD and EMM formalize interestingness using scoring functions $\varphi:2^{\mathcal{S}}\rightarrow\mathbb{R}$ (also called interestingness measures or quality measures) which map a pattern $p\in 2^{\mathcal{S}}$ to a score given a dataset $D$. A search algorithm explores the search space of patterns ($2^{\mathcal{S}}$) to find the ones with the highest scores. The computational challenge of subgroup discovery lies in the exponentially growing search space $2^{\mathcal{S}}$ with the number of selectors $|\mathcal{S}|$. Beyond straightforward constraints like pattern length or a minimum number of covered instances, optimistic estimates $oe_{\varphi}$ are integral for optimizing runtimes. An $oe_{\varphi}$ provides an upper bound on the score $\varphi$ of all specializations $p_{spec}$ of a subgroup $p_{gen}$ [lemmerich_novel_2014]: $\forall p_{spec}\supset p_{gen}:\varphi(p_{spec})\leq oe_{\varphi}(p_{gen})$. For top-k search, this can be used to disregard all specializations of $p_{gen}$ if $oe_{\varphi}(p_{gen})$ is lower than the score of all current top-k patterns. Furthermore, tight optimistic estimates [grosskreutz_tight_2008] provide tighter upper bounds by guaranteeing that the upper bound can be reached by a hypothetical specialization $p_{spec}$, for an arbitrary subset of instances $I\subseteq sg(p_{gen})$ covered by $p_{gen}$. Optimistic estimates can be integrated into various search algorithms. However, they have to be derived for each scoring function individually.
@@ 2.3 SCaPE Model Class
For EMM, the Soft Classifier Performance Evaluation (SCaPE) model class has been proposed to discover subgroups with exceptional performance of soft classifiers [duivesteijn_understanding_2014]. It integrates the ground truth labels $Y$ and the model predictions $\hat{Y}$.
The original scoring function Relative Average Subranking Loss (RASL), written $\varphi^{rasl}$ here, proposed for SCaPE, measures the number of violations of the ranking induced by the predictions $\hat{Y}(c)$ of the given soft classifier against the ranking based on the ground truth labels $Y(c)$ within a subgroup $sg$ with pattern $p$: For each positive instance $c\in\mathcal{P}_{p}$ for a pattern $p$, it counts how many negative instances $c^{\prime}\in\mathcal{N}_{p}$ are ranked higher. This count is denoted as $PEN(c,p)$. The Average Ranking Loss (ARL) is then defined as the average of $PEN(c,p)$ across all positive instances. The final scoring function $\varphi^{rasl}$ is defined relative to the overall dataset: $\varphi^{rasl}(p):=ARL(p)-ARL(\mathcal{\emptyset})$.
Although the ARL score is normalized by the number of positive instances, it is heavily influenced by the number of overall instances covered by a pattern as the penalty score $PEN(c,p)$ counts the absolute number of violations. This favors large subgroups or subgroups with heavily biased class distributions without an option to manually trade off these aspects (see Section 4.4 for details).
@@ 3 The SubROC Framework
The SubROC framework consists of three components: defining novel scoring functions based on the widely employed performance measures (ROC AUC and Precision-Recall (PR) AUC) while accounting for class imbalance, size, and redundancy (Section 3.1), deriving optimistic estimates for efficient search space exploration and pruning (Section 3.2), and significance filtering for increasing the robustness and generalizability of the discovered subgroups (Section 3.3).
@@ 3.1.1 AUC-based subgroup interestingness
To guarantee well-defined scoring, we define constraints: at least one positive instance for ARL and PR AUC and at least one instance of both classes for ROC AUC. For PR AUC, we employ the widely used approximation via a linearly interpolated PR curve, with no practical consequences in our initial tests, cf. [davis_relationship_2006].
Subgroup scoring functions commonly include a factor for trading off size against the score to prevent the discovery of only small, overfitted subgroups (e.g., see [klosgen_explora_1996, leman_exceptional_2008]). We apply the same concept to our AUC-based scoring functions $\varphi^{\prime}$ and derive size-sensitive scoring functions $\varphi_{\alpha}$ as $\varphi_{\alpha}(p)=|sg(p)|^{\alpha}\cdot\varphi^{\prime}(p)$. Here, $\alpha\in\mathbb{R}_{\geq 0}$ is the trade-off parameter, where $\alpha=0$ removes the size influence.
@@ 3.1.2 Class Balance Factor
The performance measures ROC AUC and PR AUC are known to be sensitive to the Negative Class Ratio (NCR) [cook_when_2020, boyd_unachievable_2012]. This can lead to subgroups with high scores just due to high or low class imbalances rather than exceptional model performance. To counterbalance this, we introduce a user-weighted class balance factor $cb(I)$ for our scoring functions [canbek_binary_2017, datta_multiobjective_2019]. Given a set of instances $I$, it provides a maximum value close to 1 for perfectly balanced classes and lower values for more imbalanced data. We define $cb(I)$ as 00 if only one class occurs and as $\min\{\frac{|\mathcal{N}_{I}|}{|\mathcal{P}_{I}|},\frac{|\mathcal{P}_{I}|}{|\mathcal{N}_{I}|}\}$ otherwise. We combine $cb(I)$ with a scoring function $\varphi^{\prime}$ as an additional factor, forming a class balance weighted scoring function $\varphi_{\beta}(p)$ for a pattern $p$: $\varphi_{\beta}(p)=cb(sg(p))^{\beta}\cdot\varphi^{\prime}(p)$ with $\beta\in\mathbb{R}_{\geq 0}$. Scoring functions that include both the cover size weight with parameter $\alpha$ and class balance weight with parameter $\beta$ are denoted as $\varphi_{\alpha,\beta}$. This follows the intuition that ROC AUC can be deceptive for imbalanced data, and thus, balanced subgroups should be preferred [saito2015precision]. Alternatives such as the difference in class-ratios between the subgroup and the dataset could also be considered.

### v2

# SubROC: AUC-Based Discovery of Exceptional Subgroup Performance for Binary Classifiers

@@ Abstract
Machine learning (ML) is increasingly employed in real-world applications like medicine or economics, thus, potentially affecting large populations. However, ML models often do not perform homogeneously, leading to underperformance or, conversely, unusually high performance in certain subgroups (e.g., sex=female $\wedge$ marital_status=married). Identifying such subgroups can support practical decisions on which subpopulation a model is safe to deploy or where more training data is required. However, an efficient and coherent framework for effective search is missing. Consequently, we introduce SubROC, an open-source, easy-to-use framework based on Exceptional Model Mining for reliably and efficiently finding strengths and weaknesses of classification models in the form of interpretable population subgroups. SubROC incorporates common evaluation measures (ROC and PR AUC), efficient search space pruning for fast exhaustive subgroup search, control for class imbalance, adjustment for redundant patterns, and significance testing. We illustrate the practical benefits of SubROC in case studies as well as in comparative analyses across multiple datasets.
@@ 1 Introduction
[…]
@@
Figure 1: Illustrative example. Predictions of a binary soft classifier produce a ranking of patients, from healthy to sick. The clearer this separation, the better the classifier. Here, the density distributions of healthy and diseased patients (red and blue curves, middle left) are clearly distinguishable but there is no perfect separation. SubROC automatically discovers interpretable subgroups where the model predictions create the best or worst separation of the classes (middle right). Results may highlight arbitrary sensitive subgroups of a population, where the model under- or over-performs compared to the rest of the dataset (right).
@@
Background and problem setting. The need and potential of identifying and analyzing subgroups has been widely recognized. For example, a wide array of work aims at alleviating consistent underperformance in known [2] as well as previously unknown [3] subgroups. Similar observations have been made concerning the fairness of ML models [1]. Consequently, various approaches explore such atypical subgroups systematically, e.g., via interactive tools for manual discovery [4, 5]. However, this is labor-intensive, inefficient, and poses the threat of missing essential subgroups. Thus, automated approaches enable the search for subgroups where ML models over- or underperform [6, 7]. The most principled approach employs Exceptional Model Mining (EMM) with a custom performance measure for efficient exhaustive search [8]. However, none of these approaches combines the automated and exhaustive search of EMM, standard measures for soft classifiers like ROC AUC or PR AUC, efficient search, and handling of class imbalance and statistical significance testing, which are required for a routine and practical analysis.
@@
Contribution. Consequently, we introduce SubROC, which builds on EMM to reliably and efficiently find strengths and weaknesses of classification models: We (a) propose a new group of scoring functions based on the area-under-the-curve (AUC) statistic, (b) derive tight optimistic estimates enabling fast exhaustive subgroup search, (c) provide a well-rounded framework that includes control for class imbalance, pruning of redundant results, and significance testing, (d) show the practical benefits of our approach across multiple datasets, and (e) integrate SubROC in the easy-to-use, open source pysubgroup package and provide the code for our experiments.  [^1] https://github.com/bckrlab/subroc, archival version at [9]
@@
SubROC will enable practitioners to routinely evaluate ML models by finding subgroups where they over- or underperform. This will not only improve quality control procedures and increase the confidence in deployed ML models but also lay the foundation for downstream approaches to systematically improve their performance.
@@ 2 Preliminaries
2.1 Notation
$D=(\mathcal{I},\mathcal{A})$ is a dataset with instances $c\in\mathcal{I}$ and attributes (features) $\mathcal{A}=(X_{1},\ldots,X_{m})$. $X_{i}(c)$ denotes the value of attribute $X_{i}$ in instance $c$. $Y(c)\in\{1,0\}$ is the ground truth label and $\hat{Y}(c)\in[0,1]$ a continuous (soft) prediction. We call instances with $Y(c)=1$ positives ($\mathcal{P}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=1\}$) and the rest negatives ($\mathcal{N}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=0\}$). A classifier performs well if it returns high (low) values for positives $\mathcal{P}_{\mathcal{I}}$ (negatives $\mathcal{N}_{\mathcal{I}}$).
@@ 2.2 Soft Classifier Evaluation
Traditional subgroup discovery on $Y\neq\hat{Y}$ is applicable to hard classification but ignores classifier confidence. Since in practice most models provide probability estimates, SubROC focuses on soft classification with the performance measures Receiver Operating Characteristic (ROC) area under the curve (AUC) and Precision-Recall (PR) AUC that require labels $Y$ as well as prediction scores $\hat{Y}$. This leads to conceptually different, not necessarily better results.
@@
While many evaluation measures for soft classification exist, including binary cross-entropy or the Brier score, the AUC of the ROC curve is among the most widely used. It assumes a ranking of instances according to their predicted probabilities from high to low. It can be interpreted as the probability of a randomly chosen positive instance being ranked above a randomly chosen negative instance.
@@
One reason for the popularity of AUC-based performance measures is that they summarize classifier performance across all potential thresholds that a user can apply on soft classifier probabilities to separate samples into two classes e.g., depending on misclassification costs. This makes the ROC AUC a natural choice for interpretable subgroup descriptions, contributing to the insights our framework may provide to experts. Nevertheless, ROC AUC has also been criticized, for example, for being misleading under strong class imbalance and for treating all thresholds as equally relevant.
2.3 Subgroup Discovery / Exceptional Model Mining
@@
A selector $s\in\mathcal{S}$ is a Boolean function based on a single attribute, e.g., $sex=female$ or $age\in[20,30)$. A pattern $p\in 2^{\mathcal{S}}$ is a conjunction of selectors, e.g., $sex=female\wedge age\in[20,30)$. Given a dataset, a pattern’s cover $sg(p)$ represents the set of all instances covered by it, i.e., all instances with $p=true$. A pattern and its cover are referred to as a subgroup. A subgroup $p_{spec}$ is a specialization of another subgroup $p_{gen}$ (its generalization) if its set of selectors is a proper superset of those of the other subgroup: $p_{gen}\subset p_{spec}$.
@@
Subgroup discovery (SD) [10] finds subgroups that show an interesting deviation with respect to a target concept, which is traditionally given as a single data attribute. E.g., an interesting subgroup may contain a higher-than-average disease occurrence. EMM [11] extends target concepts to model classes. Here, a subgroup is considered interesting if the model parameters fitted to the subgroup are significantly different from those fitted to the complete dataset.
@@
SD and EMM formalize interestingness using scoring functions $\varphi:2^{\mathcal{S}}\rightarrow\mathbb{R}$ (also called interestingness measures or quality measures) which map a pattern $p\in 2^{\mathcal{S}}$ to a score, given a dataset $D$. A search algorithm explores the search space of patterns ($2^{\mathcal{S}}$) to find the ones with the highest scores. The computational challenge of subgroup discovery lies in the exponentially growing search space $2^{\mathcal{S}}$ with the number of selectors $|\mathcal{S}|$. Beyond straightforward constraints like the number of selectors in a pattern or the minimum support, in terms of the number of covered instances, optimistic estimates are integral for optimizing runtimes. An optimistic estimate on a score $\varphi$}, $oe_{\varphi}$ provides an upper bound on the score $\varphi$ of all specializations $p_{spec}$ of a subgroup $p_{gen}$ [12]: $\forall p_{spec}\supset p_{gen}:\varphi(p_{spec})\leq oe_{\varphi}(p_{gen})$. When searching for the top-k patterns with the highest interestingness, this can be used to prune all specializations of $p_{gen}$ if $oe_{\varphi}(p_{gen})$ is lower than the score of all current top-k patterns. Furthermore, tight optimistic estimates [13] provide tighter upper bounds by guaranteeing that the upper bound can be reached by a hypothetical specialization $p_{spec}$, for an arbitrary subset of instances $I\subseteq sg(p_{gen})$ covered by $p_{gen}$. Optimistic estimates can be integrated into various search algorithms. However, they have to be derived for each scoring function individually.
@@ 2.4 SCaPE Model Class
For EMM, the Soft Classifier Performance Evaluation (SCaPE) model class has been proposed to discover subgroups with exceptional performance of soft classifiers [8]. It integrates the ground truth labels $Y$ and the model predictions $\hat{Y}$.
The original scoring function Relative Average Subranking Loss (RASL) $\varphi^{rasl}$ proposed for SCaPE, measures the number of violations of the ranking induced by the predictions $\hat{Y}(c)$ of the given soft classifier, against the ranking based on the ground truth labels $Y(c)$: For each positive instance $c\in\mathcal{P}_{p}$ for a pattern $p$, it counts how many negative instances $c^{\prime}\in\mathcal{N}_{p}$ are ranked higher. This count is denoted as $PEN(c,p)$. The Average Ranking Loss (ARL) is then defined as the average of $PEN(c,p)$ across all positive instances. The final scoring function $\varphi^{rasl}$ is defined relative to the overall dataset: $\varphi^{rasl}(p):=ARL(p)-ARL(\mathcal{\emptyset})$, where the empty pattern $\emptyset$ corresponds to the subgroup containing all instances.
Although the ARL score is normalized by the number of positive instances, it is heavily influenced by the number of overall instances (of either class) covered by a pattern as the penalty score $PEN(c,p)$ counts the absolute number of violations. This favors large subgroups or heavily biased class distributions without an option to manually trade off these aspects (see Section˜4.4 for details).
In practice, ARL is not a common score to measure classifier performance and the corresponding RASL measure [8] does not provide optimistic estimates for efficient subgroup search. ROC and PR AUC, which are widely applied to evaluate soft classifiers, currently are not available for the SCaPE model class. However, due to their widespread adoption, they are more readily interpretable and have technical advantages like their independence of the number of evaluated samples (also see Section˜4.4). Thus, SubROC introduces ROC and PR AUC for the SCaPE model class and provides optimistic estimates for ROC and PR AUC as well as RASL.
@@ 3 The SubROC Framework
The SubROC framework consists of three components: defining novel scoring functions based on ROC AUC, while accounting for class imbalance, size, and redundancy (Section˜3.1), deriving optimistic estimates for efficient search space exploration and pruning (Section˜3.2), and significance filtering for increasing the robustness and generalizability of the discovered subgroups (Section˜3.3). The same construction is also applied to PR AUC by replacing ROC ’s true and false positive rates with precision and recall.
@@ 3.1.1 AUC-based subgroup interestingness
To guarantee well-defined scoring, we define constraints: at least one positive instance for ARL and PR AUC and at least one instance of both classes for ROC AUC. For PR AUC, we employ the widely used approximation via a linearly interpolated PR curve, with no practical consequences in our initial tests, cf. [14].
Subgroup scoring functions commonly include a factor for trading off size against the score to prevent the discovery of only small, overfitted subgroups (e.g., see [10, 11]). We apply the same concept to our AUC-based scoring functions $\varphi^{\prime}$ and derive size-sensitive scoring functions $\varphi_{\alpha}$ as $\varphi_{\alpha}(p)=|sg(p)|^{\alpha}\cdot\varphi^{\prime}(p)$. Here, $\alpha\in\mathbb{R}_{\geq 0}$ is the trade-off parameter, where $\alpha=0$ removes the size influence.
@@ 3.1.2 Class Balance Factor
The performance measures ROC AUC and PR AUC are known to be sensitive to the Negative Class Ratio (NCR) [15, 16]. This can lead to subgroups with high scores just due to high class imbalances rather than exceptional model performance. To counterbalance this, we introduce a user-weighted class balance factor $cb(I)$ for our scoring functions [17, 18]. Given a set of instances $I$, it provides a maximum value close to 1 for perfectly balanced classes and lower values for more imbalanced data. We define $cb(I)$ as 0 if only one class occurs and as $\min\{\frac{|\mathcal{N}_{I}|}{|\mathcal{P}_{I}|},\frac{|\mathcal{P}_{I}|}{|\mathcal{N}_{I}|}\}$ otherwise. We combine $cb(I)$ with a scoring function $\varphi^{\prime}$ as an additional factor, forming a class balance weighted scoring function $\varphi_{\beta}(p)$ for a pattern $p$: $\varphi_{\beta}(p)=cb(sg(p))^{\beta}\cdot\varphi^{\prime}(p)$ with $\beta\in\mathbb{R}_{\geq 0}$. Scoring functions that include both the cover size weight with parameter $\alpha$ and class balance weight with parameter $\beta$ are denoted as $\varphi_{\alpha,\beta}$. This follows the intuition that ROC AUC can be deceptive for imbalanced data, and thus, balanced subgroups should be preferred [19]. Alternatives such as the difference in class-ratios between the subgroup and the dataset could also be considered.